OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a18 = 97, a20 = 281

OpenStudy (anonymous):

an = -1467 - 92(n + 1) an = -1467 + 92(n + 1) an = -1467 + 92(n - 1) an = -1467 + 92(n - 2) like how do i fine ninth term and all that

OpenStudy (amoodarya):

a20-a18=a1+19d-a1-17d=2d =281-97 so d=92

OpenStudy (amoodarya):

can you go on ?

OpenStudy (anonymous):

uhh im confused

OpenStudy (amoodarya):

an=a1+(n-1)d you know that ?

OpenStudy (anonymous):

yeah, so im plugging that equation into the answer

OpenStudy (amoodarya):

a18=a1+17d =97 a20=a1+19d=281 solve the system of equation

OpenStudy (anonymous):

where are you getting the 17d and the 19d from

OpenStudy (amoodarya):

a20-a18 =2d a20 - a18=281-97 so 2d =281-97=184 d=92

OpenStudy (amoodarya):

now find a1 a18 = a1+17 (92) =97 so a1=97 - 17(92)=-1467

OpenStudy (amoodarya):

can you go on ?

OpenStudy (anonymous):

so is it D

OpenStudy (anonymous):

an = -1467 + 92(n - 2)

OpenStudy (amoodarya):

no a(n)=a1+(n-1)d so an=-1467+92(n-1)

OpenStudy (amoodarya):

understand ?

OpenStudy (anonymous):

yeah thnx